OpenStudy (anonymous):

Help Determine if the Equation is Exact and integrate if so. (2xy+3)dx+(x^2-1)dy=0

OpenStudy (anonymous):

How do I know if the equation is exact?

OpenStudy (anonymous):

You know the equation is exact if the partial derivative with respect to y of the first parentheses is the same as the partial derivative with respet to x of the second parentheses.

OpenStudy (tkhunny):

3xy + 3 You already have dx, now do dy x^2 - 1 You already have dy, now do dx

OpenStudy (anonymous):

what do you mean by that?

OpenStudy (anonymous):

And the equation is exact because d/dy of (2xy + 3) is 2x and d/dx of (x^2 - 1) is 2x. Since they're both the same, the equation is exact. Do you see what I mean?

OpenStudy (anonymous):

that makes sense thank you!

OpenStudy (anonymous):

could you please help with integrating as well?

OpenStudy (anonymous):

I'm trying to work it out, I haven't done differential equations in a while.

OpenStudy (anonymous):

i've never done differential equations, this is my first semester taking it, take your time

OpenStudy (anonymous):

The idea of this is that a function w(x,y), where y is also a function of x, has a derivative dw/dx + (dw/dy) * (dy/dx). If you divide your equation up there by dx, you get (2xy + 3) + (x^2-1)dy/dx = 0. If you compare these two, you see that dw/dx = 2xy + 3 and dw/dy = x^2-1, which works if it's an exact equation. So what this is saying is that there's a function of x and y (that we call w) whose derivative is the LHS of the equation you gave.

OpenStudy (anonymous):

dw/dx = 2xy + 3 dw = (2xy + 3)dx w = yx^2 + 3x + f(y) Where f(y) is kind of like + C, so we have to find f(y) now.

OpenStudy (tkhunny):

(2xy+3)dx+(x^2-1)dy=0 IF it is exact, we have some wonderful function F(x,y), such that: \(F_{x} = 2xy + 3\;and\;F_{y} = x^{2}-1\) We have not yet proven that it is exact, so that is the first thing to do. Demonstrate that \(F_{xy} = F_{yx}\) Here we go. Given \(F_{x} = 2xy + 3\;we\;have\;F_{xy} = 2x\) Given \(F_{y} = x^{2} - 1\;we\;have\;F_{yx} = 2x\) Since \(F_{xy} = F_{yx}\), we call it "Exact".

OpenStudy (anonymous):

also if determined exact, does that eliminate them from being separable, linear, or none of these?

OpenStudy (anonymous):

Take partial derivative with respect to y. dw/dy = x^2 + f'(y) We already know an expression for dw/dy from before so set this equal to that. x^2 + f'(y) = x^2 - 1 f'(y) = -1 f(y) = -y So w is HOPEFULLY: w = yx^2 + 3x - y

OpenStudy (tkhunny):

Why do we care? It's exact. The solution is relatively simple. It's possible some could be put into other forms, but again, why?

OpenStudy (anonymous):

thank you bagajr i also got x^2y+3x-y=C

OpenStudy (tkhunny):

Just for fun, start with the other one and see if you get the same result!

OpenStudy (anonymous):

i did on paper and it is

OpenStudy (tkhunny):

Super!